# Generating Hypotheses of Dynamic Causal Graphs in Neuroscience: Leveraging Generative Factor Models of Observed Time Series

### v1

@@ 1 Introduction
Thus, we focus on generating hypotheses about mechanisms governing nonlinear and state-dependent (or “dynamic”) behavior in time series systems (hereafter referred to as “temporal systems” or simply “systems” to conserve space). While understandable concerns may be raised about the detectability of such mechanisms in even simple systems (see our discussion on identifiability in Appendix A.1), prior literature has demonstrated that using automated hypotheses generation for these mechanisms can still yield useful insights which aid scientific discovery (Mague et al., 2022). Indeed, we accurately estimate causal graphs in observed systems without making restrictive assumptions as to the state-dependence or linearity/nonlinearity of causal relationships. We do this by combining key concepts from both factor-based and nonlinear models of causal time series.
@@ 2 Related Work
There are numerous causal discovery methods which avoid Granger causality (Runge, 2020; Gerhardus & Runge, 2020; Zou & Feng, 2009; Yu et al., 2019). These methods do not assume causality from temporal correlation, and thus are better positioned to handle causal relations such as contemporaneous effects (Runge, 2020) and latent confounders (Gerhardus & Runge, 2020). However, these methods tend to rely on assumptions that rule out hypotheses we would like to explore in neuroscience; for example, many methods assume causal graphs are acyclical (Runge, 2020; Gerhardus & Runge, 2020; Zou & Feng, 2009; Yu et al., 2019), which recurrent, multi-time scale systems like the brain may fundamentally violate - e.g. through cross-frequency coupling (Dzirasa et al., 2010; Allen et al., 2011). Thus, our approach differs from these methods and we instead report empirical comparisons between our method and Regime-PCMCI (Saggioro et al., 2020) in Section 4.2.
@@
Additionally, we note there is work using dynamic nonlinear models through time, such as State-Dependent Causal Inference (SDCI) (Rodas et al., 2021), Switching Neural Network Trackers (SNNTs) (Ghimire, 2018), and Constraint-based causal Discovery from Nonstationary/heterogeneous Data (CD-NOD) (Zhang et al., 2017a). However, SDCI models connections between embedded latent variables (which embedding process would hamper our ability to identify variables with a single physical signal/brain region) and assigns a state label to each latent variable in the system, a granularity which we cannot attain with labels from behavioral assays in neuroscience. Meanwhile, SNNTs are probabilistic in nature instead of being designed for causal discovery and use discrete states instead of the superposition of factors in the datasets used in our Section 4.3 experiments and the design of our proposed method. In contrast, the development of CD-NOD was largely motivated by the task of performing causal discovery on neurological (esp. fMRI) data, and it thus shares similarities with our work in much of the problem setup and certain assumptions. Still, there are enough differences between our approach and CD-NOD - such as our focus on autoregressive systems and unconstrained directionality of underlying causal relationships - that our direct comparisons with Regime-PCMCI (Saggioro et al., 2020) can be considered sufficient for our purposes.
@@ 3.1 Problem Statement
Suppose we must hypothesize a causal model for a dataset $\mathcal{D}$ of $N\in\mathbb{N}$ multi-dimensional (i.e. ‘multi-channel’) time series recordings $\mathbf{x}_{n}$. Let $\mathcal{D}=(\mathbf{X})=(\mathbf{x}_{n})_{n=1}^{N}$ denote the complete dataset of recordings. Each recording $\mathbf{x}_{n}\in\mathbf{X}$ contains the observed history of $\mathrm{n_{c}}$ system variables at regular intervals with $T\in\mathbb{N}$ temporal measurements. Let $\tau_{\mathrm{in}}\in\mathbb{N}_{\leq T}$ be the maximal number of time steps over which we estimate causal relationships. For simplicity, assume observations are of scalar features such that $\mathbf{x}_{n,t:t+\tau_{\mathrm{in}}}\in\mathbb{R}^{\mathrm{n_{c}}\times\tau_{\mathrm{in}}}\;\forall\;t\leq T\;\mathrm{and}\;n\leq N$.
We assume our data is a noisy realization from a time-dependent system $\Phi(t)$. We must learn a model that a) reasonably approximates $\Phi$ at any point in time and b) estimates causal influences between system variables in $\Phi$. We approximate $\Phi$ by learning a factor model $f_{\phi}$ over the set of observed variables, where the expression of each factor is dynamic over time. Indeed, in Sections 4.1, 4.2, and 4.3 we implement $\Phi$ as a factor model itself (see Section  3.2 for formalization). The functions comprising each factor can take on many forms including hypothesis generation functions. Additionally, the linear combinations in factor models enable us to track statistics such as presence of a factor over time or in relationship to auxiliary variables (e.g. behavior), making these models useful for scientific research.
@@ 3.2 General Notation
where $K^{*}$ is the number of true factors $\mathcal{F}_{k}\in\mathcal{F}$ acting on all system variables $\mathcal{V}$ at each time step $t$ i.e. $\mathcal{F}_{k}:\mathbb{R}^{\mathcal{V}}\rightarrow\mathbb{R}^{\mathcal{V}}$ according to the system state $\mathcal{G}(t)\in\mathbb{R}^{K^{*}}$.
@@
We now discuss how estimates of these relationships can be formally represented during evaluation. We turn to graph theory to express each factor’s estimate of the strength of Granger causal relationships as a series of adjacency matrices mapping past variable (i.e. ‘node’) states/values to the present. The flexibility of graphs lets us also consider an abstract adjacency matrix summarizing the effect of $\Phi(t)$ on variables in the true system. Formally, we summarize our models’ estimated Granger causal relationships in a series of adjacency matrices $\hat{\mathbf{A}}\in\mathbb{R}^{\mathrm{n_{c}}\times\mathrm{n_{c}}\times\tau_{\mathrm{in}}}$ where each $a_{i,j,t}\in\hat{\mathbf{A}}$ is the estimated weight of the Granger causal relationship where the state of node $\nu_{j}\in\mathcal{V}$ from $t$ time steps in the past ‘causally’ effects current node $\nu_{i}\in\mathcal{V}$. To make quantitative comparisons, we standardized the representation(s) of baselines’ and our models’ estimated causal graph(s) by summing over time-lagged features (see Appendix B.2); hence, we define the lag-summed view of $\hat{\mathbf{A}}$ as
@@
$\odot^{\dagger}$ ‘broadcast multiplication’ (Section 3.2)
* (emphasizes) scalar-vector multiplying
$\overline{(...)}$ Average of value of ()(...)( … )
‘MSE’ The mean-squared error penalty
@@ 3.3 Base (Single-Factor) Model
Regarding the loss function of this single-factor model, we assume no direct knowledge of the target system’s underlying causal graph since we are performing hypothesis generation. Instead, we estimate causal connections in an unsupervised fashion. Hence, our objective function is essentially a sum over regularization terms guiding the model towards solutions with desirable qualities, including sparsity.
@@
where $\eta,\omega\in\mathbb{R}$ are chosen hyperparameters. Again, the $L1$ regularization term here serves two purposes for the factor’s causal model: 1) to encourage sparsity/parsimony and 2) to emphasize reliance on more recent time lags via the log term. We use the Adam optimization algorithm to train our model to minimize the above loss function. We also include weight decay (i.e. $L2$-norm regularization) on all parameters as done in prior work (Tank et al., 2021).
@@ 3.4 RElational Discovery via ConditionaLly Interacting Forecasting Factors (REDCLIFF)
To learn factor coefficients for each factor, we introduce a ‘state model’ $g_{\theta}$ which is a parameterized function that generates factor weights (or ‘scores’) conditioned on the history of the system. Note $g_{\theta}$ can be seen as trying to classify each factor as either likely or unlikely to be needed in each time series forecast. Thus $g_{\theta}$ is fundamentally different from the generative factors of $f_{\phi}$ (which perform regression on the time series itself); indeed, we found that $g_{\theta}$ typically required more historical information than the generative factors to sufficiently learn its task (see Section 3.6 and Appendix A.2). Thus, we provide additional historical information of length $\tau_{cl}\in\mathbb{N}_{\leq T}$ to $g_{\theta}$ so $g_{\theta}$ sees a combined context window of $\tau_{\mathrm{in}}+\tau_{cl}$ time steps.
@@
The objective function for training REDCLIFF models has two major components. Modifying Equation 3, the first component $\mathcal{L}_{f}$ pertains to the generative factors $f_{\phi_{k}}$ and retains the MSE forecasting penalty, but now includes all $\mathrm{n_{k}}\geq 1$ factors in the $L1$-norm penalty. We also include a cosine similarity penalty to induce our prior that each factor should serve different purposes by encouraging factors to dissociate (details in Appendices B.3 and C.6). Specifically,
@@ 3.5 REDCLIFF-Supervised (REDCLIFF-S)
Now suppose $\mathcal{D}$ includes a set $\mathbf{Y}$ of “global state” (i.e. “behavioral state” in our motivating neurobehavioral cases) labels for the span of each recording in $\mathbf{X}$; that is, $\mathcal{D}=(\mathbf{X},\mathbf{Y})$. We assume the set $\mathbf{Y}$ takes the form $\mathbf{Y}=(\mathbf{y}_{n}\in\mathbb{R}^{B\times T})_{n=1}^{N}$ where $B\in\mathbb{N}$ is the number of global state variables and both $N$ and $T$ are as before.
We seek to estimate causal patterns linked to specific behaviors in signals. Thus, we follow prior literature and assign behaviors to the first $B$ elements of $\boldsymbol{\alpha}$ (Mague et al., 2022). Because the desired factor scores in $\boldsymbol{\alpha}_{:B}$ may not always correlate with the behaviors they are associated with, we define invertible functions $g_{\alpha}$ and $g_{y}$ such that $\hat{\mathbf{y}}=g_{y}(g_{\alpha}^{-1}(\boldsymbol{\alpha}_{:B}))$. These invertible functions provide more flexibility than directly copying elements of $\boldsymbol{\alpha}$ while still preserving the relative information contained therein, and may be used to limit the effect gradients from the forecasting loss have on $g_{\theta}$. When the behavioral label $\mathbf{y}_{n}$ is a known quantitative variable, we can generate $\hat{\mathbf{y}}_{n}$ via supervised regression. Alternatively, if $\mathbf{y}_{n}$ is categorical, we can predict a behavior as ‘present’ in the signal when the value of one of these behavioral elements crosses a threshold $c_{b}\in\mathbb{R}$.
@@ 3.6 Training Strategies and Details
We train REDCLIFF(-S) using the Adam optimizer with weight decay. To effectively learn each model, we first pretrain the state model $g_{\theta}$ using the $\mathcal{L}_{g}$ and $\lambda\mathrm{MSE}(\mathbf{Y},\hat{\mathbf{Y}})$ terms from Equation 9 (setting $\lambda=0$ in the unsupervised case). Then, we freeze the state model and train the generative factors in $f_{\phi}$ using Equation 5 (we call this “acclimation”). Following these steps, we employ the full training objective (Equation 7 or 9) to jointly update all parameters.
One challenge for hypothesis generation is that you cannot perform model selection explicitly on holdout graph discovery performance on real data as the true graph is unknown. Instead, we stop training once the number of maximal iterations is reached or when the following criterion is minimized on the validation set $\mathcal{D}_{\mathrm{v}}=(\mathbf{X}_{\mathrm{v}},\mathbf{Y}_{\mathrm{v}})$:
@@
In terms of computational complexity, the REDCLIFF(-S) implementations we tested in this work scale (in training and testing) as the product of the number of factors $\mathrm{n_{k}}$ and the number of forward passes $n_{\mathrm{sim}}$ made by each factor  [^2] We experimented with having $n_{\mathrm{sim}}\in\mathbb{N}_{\geq 1}$ forward passes before updating REDCLIFF-S parameters in some grid searches. In practice, $n_{\mathrm{sim}}=1$ performed best, but we include the general form in our complexity analyses for rigor. plus the computational complexity of the state model. More formally, let $C_{\phi_{*}}$ represent the maximal (temporal or spacial, depending on context) complexity of any factor in $\phi$ and $C_{\theta}$ be the complexity of the state model; then the temporal complexity of our implementation is $\mathcal{O}(\mathrm{n_{k}}\cdot n_{\mathrm{sim}}\cdot C_{\phi_{*}}+C_{\theta})$ and the spacial complexity is the same. Learning can also be parallelized over factors for scalable implementations.
@@ 4 Experiments on Simulated Datasets
We now report empirical comparisons of REDCLIFF-S against other methods [^3] Code has been made available at github.com/carlson-lab/redcliff-s-hypothesizing-dynamic-causal-graphs, with additional results given in Appendix C. All algorithms were trained on a local cluster featuring a variety of GPU devices, including A6000 GPUs.
We focus much of our analysis on f1-score performance, partly because our neuroscientific application implicitly assumes all parts of the brain are causally related to some degree and the difficulty is determining which connections are the most relevant at key moments; the f1-score’s emphasis on the presence (positive label) of relationships is thus well suited to our setting. As is common practice, we report ‘optimal’ f1-scores in which we compute each evaluated algorithm’s f1-score using a classification threshold (i.e. the value at which logits are labeled negative or positive) which yields the highest possible f1-score for that algorithm on the task. We do use the “Adjustment Identification Distance” (or AID) family of causal distance metrics (Henckel et al., ) in our comparisons against several supervised baselines (Table 2 and Appendix C.3), since all algorithms in this comparison had access to ‘global state’ information that relates to causal edge relevancy, possibly making it harder for the f1-score to differentiate algorithms. We also report some ROC-AUC scores, such as in Table 2 and the Appendices.
Figure 2: Synthetic Systems results from select systems. A) Average optimal f1-score and standard error of the mean (SEM) between true and estimated inter-variable causal relationships. B) Average pair-wise improvement and SEM between optimal f1-scores obtained by REDCLIFF-S and baselines.

### v2

@@ 1 Introduction
Thus, we focus on generating hypotheses about mechanisms governing nonlinear and state-dependent (or “dynamic”) behavior in time series systems (hereafter referred to as “temporal systems” or simply “systems” to conserve space). While understandable concerns may be raised about the detectability of such mechanisms in even simple systems (see our discussion on identifiability in Appendix A.1), prior literature has demonstrated that using automated hypothesis generation for these mechanisms can still yield useful insights which aid scientific discovery (Mague et al., 2022). Indeed, we accurately estimate causal graphs in observed systems without making restrictive assumptions as to the state-dependence or linearity/nonlinearity of causal relationships. We do this by combining key concepts from both factor-based and nonlinear models of causal time series.
@@ 2 Related Work
There are numerous causal discovery methods which avoid Granger causality (Runge, 2020; Gerhardus & Runge, 2020; Zou & Feng, 2009; Yu et al., 2019). These methods do not assume causality from temporal correlation, and thus are better positioned to handle causal relations such as contemporaneous effects (Runge, 2020) and latent confounders (Gerhardus & Runge, 2020). However, these methods tend to rely on assumptions that rule out hypotheses we would like to explore in neuroscience; for example, many methods assume causal graphs are acyclical (Runge, 2020; Gerhardus & Runge, 2020; Zou & Feng, 2009; Yu et al., 2019), which recurrent, multi-time scale systems like the brain may fundamentally violate - e.g., through cross-frequency coupling (Dzirasa et al., 2010; Allen et al., 2011). Thus, our approach differs from these methods and we instead report empirical comparisons between our method and Regime-PCMCI (Saggioro et al., 2020) in Section 4.2.
@@
There is also work using dynamic nonlinear models through time, such as State-Dependent Causal Inference (SDCI) (Rodas et al., 2021), Switching Neural Network Trackers (SNNTs) (Ghimire, 2018), and Constraint-based causal Discovery from Nonstationary/heterogeneous Data (CD-NOD) (Zhang et al., 2017a). However, SDCI models connections between embedded latent variables, which would hamper identifying variables with a single physical signal/brain region, and assigns a state label to each latent variable in the system, a granularity which we cannot attain with labels from behavioral assays in neuroscience. Meanwhile, SNNTs are probabilistic in nature instead of being designed for causal discovery and use discrete states instead of a superposition of factors used in our Section 4.3 experiments and the design of our proposed method. In contrast, the development of CD-NOD was largely motivated by the task of performing causal discovery on neurological (esp. fMRI) data, and it thus shares similarities with our work in much of the problem setup and certain assumptions. Still, there are enough differences between our approach and CD-NOD - such as our focus on autoregressive systems and unconstrained directionality of underlying causal relationships - that our direct comparisons with Regime-PCMCI (Saggioro et al., 2020) can be considered sufficient for our purposes.
@@ 3.1 Problem Statement
Suppose we must hypothesize a causal model for a dataset $\mathcal{D}$ of $N\in\mathbb{N}$ multi-dimensional (i.e., ‘multi-channel’) time series recordings $\mathbf{x}_{n}$. Let $\mathcal{D}=(\mathbf{X})=(\mathbf{x}_{n})_{n=1}^{N}$ denote the complete dataset of recordings. Each recording $\mathbf{x}_{n}\in\mathbf{X}$ contains the observed history of $\mathrm{n_{c}}$ system variables at regular intervals with $T\in\mathbb{N}$ temporal measurements. Let $\tau_{\mathrm{in}}\in\mathbb{N}_{\leq T}$ be the maximal number of time steps over which we estimate causal relationships. For simplicity, assume observations are of scalar features such that $\mathbf{x}_{n,t:t+\tau_{\mathrm{in}}}\in\mathbb{R}^{\mathrm{n_{c}}\times\tau_{\mathrm{in}}}\;\forall\;t\leq T\;\mathrm{and}\;n\leq N$.
We assume our data is a noisy realization from a time-dependent system $\Phi(t)$. We must learn a model that a) reasonably approximates $\Phi$ at any point in time and b) estimates causal influences between system variables in $\Phi$. We approximate $\Phi$ by learning a factor model $f_{\phi}$ over the set of observed variables, where the expression of each factor is dynamic over time. Indeed, in Sections 4.1, 4.2, and 4.3 we implement $\Phi$ as a factor model itself (see Section  3.2 for formalization). The functions comprising each factor can take on many forms including hypothesis generation functions. Additionally, the linear combinations in factor models enable us to track statistics such as presence of a factor over time or in relationship to auxiliary variables (e.g., behavior), making these models useful for scientific research.
@@ 3.2 General Notation
where $K^{*}$ is the number of true factors $\mathcal{F}_{k}\in\mathcal{F}$ acting on all system variables $\mathcal{V}$ at each time step $t$ i.e., $\mathcal{F}_{k}:\mathbb{R}^{\mathcal{V}}\rightarrow\mathbb{R}^{\mathcal{V}}$ according to the system state $\mathcal{G}(t)\in\mathbb{R}^{K^{*}}$.
@@
We now discuss how estimates of these relationships can be formally represented during evaluation. We turn to graph theory to express each factor’s estimate of the strength of Granger causal relationships as a series of adjacency matrices mapping past variable (i.e., ‘node’) states/values to the present. The flexibility of graphs lets us also consider an abstract adjacency matrix summarizing the effect of $\Phi(t)$ on variables in the true system. Formally, we summarize our models’ estimated Granger causal relationships in a series of adjacency matrices $\hat{\mathbf{A}}\in\mathbb{R}^{\mathrm{n_{c}}\times\mathrm{n_{c}}\times\tau_{\mathrm{in}}}$, where each $a_{i,j,t}\in\hat{\mathbf{A}}$ is the estimated weight of the Granger causal relationship where the state of node $\nu_{j}\in\mathcal{V}$ from $t$ time steps in the past ‘causally’ effects current node $\nu_{i}\in\mathcal{V}$. To make quantitative comparisons, we standardized the representation(s) of baselines’ and our models’ estimated causal graph(s) by summing over time-lagged features (see Appendix B.2); hence, we define the lag-summed view of $\hat{\mathbf{A}}$ as
@@
$\odot^{\dagger}$ ‘broadcast multiplication’ (Sec. 3.2)
* (emph.) scalar-vector multiplying
$\overline{(...)}$ Average value of ()(...)( … )
‘MSE’ The mean squared error penalty
@@ 3.3 Base (Single-Factor) Model
Regarding the loss function of this single-factor model, we assume no direct knowledge of the target system’s underlying causal graph, since we are performing hypothesis generation. Instead, we estimate causal connections in an unsupervised fashion. Hence, our objective function is essentially a sum over regularization terms guiding the model towards solutions with desirable qualities, like sparsity.
@@
where $\eta,\omega\in\mathbb{R}$ are chosen hyperparameters. Again, the $L1$ regularization term here serves two purposes for the factor’s causal model: 1) to encourage sparsity/parsimony, and 2) to emphasize reliance on more recent time lags via the log term. We use the Adam optimization algorithm to train our model to minimize the above loss function. We also include weight decay (i.e., $L2$-norm regularization) on all parameters as done in prior work (Tank et al., 2021).
@@ 3.4 RElational Discovery via ConditionaLly Interacting Forecasting Factors (REDCLIFF)
To learn factor coefficients for each factor, we introduce a ‘state model’ $g_{\theta}$ which is a parameterized function that generates factor weights (or ‘scores’) conditioned on the history of the system. Note $g_{\theta}$ can be seen as trying to classify each factor as either likely or unlikely to be needed in each time series forecast. Thus, $g_{\theta}$ is fundamentally different from the generative factors of $f_{\phi}$ (which perform regression on the time series itself); indeed, we found that $g_{\theta}$ typically required more historical information than the generative factors to effectively learn their tasks (see Section 3.6 and Appendix A.2). Thus, we provide additional historical information of length $\tau_{cl}\in\mathbb{N}_{\leq T}$ to $g_{\theta}$ so $g_{\theta}$ sees a combined context window of $\tau_{\mathrm{in}}+\tau_{cl}$ time steps.
@@
The objective function for training REDCLIFF models has two main components. Modifying Equation 3, the first component $\mathcal{L}_{f}$ pertains to the generative factors $f_{\phi_{k}}$ and retains the MSE forecasting penalty, but now includes all $\mathrm{n_{k}}\geq 1$ factors in the $L1$-norm penalty. We also include a cosine similarity penalty to induce our prior that each factor should serve different purposes by encouraging factors to dissociate (details in Appendices B.3 and C.6). Specifically,
@@ 3.5 REDCLIFF-Supervised (REDCLIFF-S)
Now suppose $\mathcal{D}$ includes a set $\mathbf{Y}$ of “global state” (i.e., “behavioral state” in our motivating neurobehavioral cases) labels for the span of each recording in $\mathbf{X}$; that is, $\mathcal{D}=(\mathbf{X},\mathbf{Y})$. We assume the set $\mathbf{Y}$ takes the form $\mathbf{Y}=(\mathbf{y}_{n}\in\mathbb{R}^{B\times T})_{n=1}^{N}$ where $B\in\mathbb{N}$ is the number of global state variables and both $N$ and $T$ are as before.
We seek to estimate causal patterns linked to specific behaviors in signals. Thus, we follow prior literature and assign behaviors to the first $B$ elements of $\boldsymbol{\alpha}$ (Mague et al., 2022). To enforce better separation between the factors $f_{\phi_{k}}$ and state model $g_{\theta}$ and because our behavioral labels may be noisy, we define invertible functions $g_{\alpha}$ and $g_{y}$ such that $\hat{\mathbf{y}}=g_{y}(g_{\alpha}^{-1}(\boldsymbol{\alpha}_{:B}))$. Invertibility is necessary for fully preserving information between these operations, allowing us to deduce the relationship between label predictions and factor weights; see discussion in Appendix A.2 for further details. When the behavioral label $\mathbf{y}_{n}$ is a known quantitative variable, we can generate $\hat{\mathbf{y}}_{n}$ via supervised regression. Alternatively, if $\mathbf{y}_{n}$ is categorical, we can predict a behavior as ‘present’ in the signal when the value of one of these behavioral elements crosses a threshold $c_{b}\in\mathbb{R}$.
@@ 3.6 Training Strategies and Details
We train REDCLIFF(-S) using the Adam optimizer with weight decay. To effectively learn each model, we first pre-train the state model $g_{\theta}$ using the terms $\mathcal{L}_{g}$ and $\lambda\mathrm{MSE}(\mathbf{Y},\hat{\mathbf{Y}})$ from Equation 9 (setting $\lambda=0$ in the unsupervised case). Then, we freeze the state model and train the generative factors in $f_{\phi}$ using Equation 5 (we call this “acclimation”). Following these steps, we employ the full training objective (Equation 7 or 9) to jointly update all parameters.
One challenge for hypothesis generation is that you cannot perform model selection explicitly using holdout graph discovery performance on real data as the true graph is unknown. Instead, we stop training once the number of maximal iterations is reached or when the following criterion is minimized on the validation set $\mathcal{D}_{\mathrm{v}}=(\mathbf{X}_{\mathrm{v}},\mathbf{Y}_{\mathrm{v}})$:
@@
In terms of computational complexity, the REDCLIFF(-S) implementations we tested in this work scale (in training and testing) as the product of the number of factors $\mathrm{n_{k}}$ and the number of forward passes $n_{\mathrm{sim}}$ made by each factor  [^2] We experimented with having $n_{\mathrm{sim}}\in\mathbb{N}_{\geq 1}$ forward passes before updating REDCLIFF-S parameters in some grid searches. In practice, $n_{\mathrm{sim}}=1$ performed best, but we include the general form in our complexity analyses for rigor. plus the computational complexity of the state model. More formally, let $C_{\phi_{*}}$ represent the maximal (temporal or spatial, depending on context) complexity of any factor in $\phi$ and $C_{\theta}$ be the complexity of the state model; then the temporal complexity of our implementation is $\mathcal{O}(\mathrm{n_{k}}\cdot n_{\mathrm{sim}}\cdot C_{\phi_{*}}+C_{\theta})$ and the spatial complexity is the same. Learning can also be parallelized over factors for scalable implementations.
@@ 4 Experiments on Simulated Datasets
We now report empirical comparisons of REDCLIFF-S against other methods [^3] Code has been made available at github.com/carlson-lab/redcliff-s-hypothesizing-dynamic-causal-graphs, with additional results given in Appendix C. All algorithms were trained using a local cluster that featured various GPU devices, including A6000 GPUs.
We focus much of our analysis on f1-score performance, partly because our neuroscientific application implicitly assumes all parts of the brain are causally related to some degree and the difficulty is determining which connections are the most relevant at key moments; the f1-score’s emphasis on the presence of positive-label relationships is thus well suited to our setting. As is common practice, we report ‘optimal’ f1-scores in which we compute each evaluated algorithm’s f1-score using a classification threshold (i.e., the value at which logits are labeled negative or positive) which yields the highest possible f1-score for that algorithm on the task. We do use the “Adjustment Identification Distance” (or AID) family of causal distance metrics (Henckel et al., ) in our comparisons against several supervised baselines (Table 2 and Appendix C.3), since all algorithms in this comparison had access to ‘global state’ information related to causal edge relevancy, possibly making it harder for the f1-score to differentiate algorithms. We also report some ROC-AUC scores, as in Table 2 and the appendices.
Figure 2: Synthetic Systems results from select systems. A) Average optimal f1-score and standard error of the mean (SEM) between true and estimated inter-variable causal relationships. B) Average pairwise improvement and SEM between optimal f1-scores obtained by REDCLIFF-S and baselines.